OpenStudy (anonymous):

The current in a bayou near lafayette, louisiana, is 4 miles per hour. A paddle boat travels 48 miles upstream in the same amount of time it tkaes to travel 72 miles downstream. Find the speed of the boat in still water? how do you set this problem up ?

OpenStudy (anonymous):

72=4t+rt 48=rt-4t

OpenStudy (anonymous):

How do i even begin to solve this problem?

OpenStudy (anonymous):

solve for t. substitute into other equation

OpenStudy (anonymous):

48=t(r-4)

OpenStudy (anonymous):

I'm so confused to be honest. :(

OpenStudy (anonymous):

so solve for t. 48/(r-4)=t

OpenStudy (anonymous):

plug it into the other equation

OpenStudy (anonymous):

72=t(4+r)

OpenStudy (anonymous):

t = d/r Let r be the speed of the boat in still water. The elapsed time for both trips is the same. The LHS is the expression for the upstream trip. Solve the following for r:\[\frac{48}{r-4}= \frac{72}{r+4}\]r = 20 mph